OpenStudy (anonymous):

isolate the radical √5x-1=x-(?

OpenStudy (anonymous):

what in the world is on the right hand side?

OpenStudy (anonymous):

thats what it wants me to find out

OpenStudy (anonymous):

\[\sqrt{5x-1}=x\]?

OpenStudy (anonymous):

then -(

OpenStudy (anonymous):

\[\sqrt{5x-1}=x-something\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

sorry i cannot imagine what the something is. do you have choices? have you seen a problem like this?

OpenStudy (anonymous):

nope hold on

OpenStudy (anonymous):

the answer is 3

OpenStudy (anonymous):

how is \[sqrt{5x-1}=x-3?

OpenStudy (anonymous):

oooooooooooooooh i understand the question. it is "solve for x"

OpenStudy (anonymous):

no that is not right either. if \[\sqrt{5x-1}=x\] then x is not 3. i don't understand what this question is asking, do you?

OpenStudy (anonymous):

it told me it was 3 but just to isolate the radical